OpenStudy (anonymous):

Could someone help me with this one...? Billy wants to factor x2 + 8x + 16 by grouping; however, Sara says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

hartnn (hartnn):

whats your attempt ?

hartnn (hartnn):

one way (of grouping) is that you first find 2 numbers whose sum is 8 and product is 16

hartnn (hartnn):

can u find these 2 numbers?

hartnn (hartnn):

yes1 so split 8x as 4x and 4x now here comes the grouping part (x^2+4x) +(4x+16)

hartnn (hartnn):

you still haven't factored right ? factor out 'x' from 1st group and 4 from 2nd group , what u get?

hartnn (hartnn):

what seems to be common from x^2+4x ? x^2=x*x and 4x=4*x

hartnn (hartnn):

i just see x common in both. so, x^2+4x = x (x+4) agree ?

hartnn (hartnn):

4x +16 = ...?

hartnn (hartnn):

yeah

hartnn (hartnn):

x is the gcf of 1st group only...

hartnn (hartnn):

whats common in 4x +16 ??

hartnn (hartnn):

the 1st group (x^2+4x) = x(x+4) 2nd group (4x+16) = 4(x+4) now whats common in both groups ?

hartnn (hartnn):

i don't see 'awesome' in any of the groups, how can it be common :P

hartnn (hartnn):

yes, 4 is gcf of 2nd group.

hartnn (hartnn):

the 1st group (x^2+4x) = x\( \huge (x+4)\) 2nd group (4x+16) = 4\( \huge (x+4) \) once again, whats common in both groups ?

hartnn (hartnn):

we r not yet done with factor by grouping...

hartnn (hartnn):

you note that x+4 is common from both group, so we can factor it out

hartnn (hartnn):